OpenStudy (mellamokatie):

Which ordered pair is a solution of the inequality? y (>=) 4x - 5 A. (3, 4) B. (2, 1) C. (3, 0) D. (1, 1)

jimthompson5910 (jim_thompson5910):

Plug in each ordered pair to test

OpenStudy (mellamokatie):

oh okay one moment

jimthompson5910 (jim_thompson5910):

For example, let's say you wanted to test the ordered pair (10,5) x = 10 y = 5 Replace each x with 10, each y with 5 \[\Large y \ge 4x - 5\] \[\Large 5 \ge 4*10 - 5\] \[\Large 5 \ge 40 - 5\] \[\Large 5 \ge 35\] which is false. So (10,5) is NOT in the solution region. This means (10,5) is NOT a solution to the inequality.

OpenStudy (mellamokatie):

Okay so so far I am on A so would it look like 4 (> =) 4(3) - 5

OpenStudy (mellamokatie):

A is not the one it looks like

jimthompson5910 (jim_thompson5910):

no need for parenthesis around the inequality sign you can say 4 >= 4(3) - 5

jimthompson5910 (jim_thompson5910):

4 >= 4(3) - 5 will turn into 4 >= 7 which is false so yes, A is not part of the solution set

OpenStudy (mellamokatie):

Oh okay

OpenStudy (mellamokatie):

So the next one is 1 >= 4(2) - 5 which is B is not correct because I get 1>=3

jimthompson5910 (jim_thompson5910):

good, so B is out too

OpenStudy (mellamokatie):

Okay so let me find out C

OpenStudy (mellamokatie):

So after doing C I get 0 >= 7 which is not right so then it must be D let me check

OpenStudy (mellamokatie):

I think its D is it? I got 1 >= 1

jimthompson5910 (jim_thompson5910):

y >= 4x - 5 y >= 4*1 - 5 ... replace x with 1 1 >= 4*1 - 5 ... replace y with 1 1 >= 4 - 5 1 >= -1 .... true

jimthompson5910 (jim_thompson5910):

Since 1 >= -1 is true, this means D is the only point in the solution set

OpenStudy (mellamokatie):

I had a feeling it was -1 I was not too sure though haha but actually it makes sense I think because 1 would be greater

jimthompson5910 (jim_thompson5910):

So as we can see on the graph (see attached image) the blue shaded region represents the solution set of \(\Large y \ge 4x-5\) The only point in that solution set is D(1,1). The other points are NOT in the blue shaded region

OpenStudy (mellamokatie):

Oh okay so it is D?

jimthompson5910 (jim_thompson5910):

yes D is the only answer

OpenStudy (mellamokatie):

Okay I have a similar one I will re post and then put the answer I think it is and tag you :D. Thanks for the medal :DD